### v1

# Design of a Wearable Parallel Electrical Impedance Imaging System for Healthcare

Abstract
A wireless wearable electrical impedance tomography (EIT) system has been developed using the AD5933 for real-time lung respiration imaging. The system uses a current injection method tailored to the human body’s impedance characteristics. It applies a voltage excitation and measures the resulting current as the voltage passes through the body. Additionally, measures are taken to suppress the effects of parasitic capacitance, which can lead to signal oscillations and leakage currents. Additionally, to improve data acquisition speed, five AD5933 units are used for parallel measurements, with multiple measures taken to ensure high synchronization during the parallel acquisition process. The results demonstrate conductivity images generated from the EIT system, with data collected from a water tank, human lungs during respiration, and a human calf at rest, confirming that this portable EIT system can measure biological tissues with high precision and low cost.
{IEEEkeywords} Electrical Impedance Tomography, Parallel Architecture, AD5933, Portable Measurement System
1 Introduction
[…]
Electrical impedance tomography (EIT) [1] is a non-invasive imaging technique in which a low-amplitude, safe alternating current is injected into the target region through electrode pairs, while the resulting boundary voltages induced by the current are measured at the remaining electrodes. The internal impedance distribution of the target region is then reconstructed using inverse-problem algorithms.
Figure 1: (a)Overview of the EIT System Architecture. (b)Photograph of the EIT System.
There are primarily two methods for measuring boundary voltages. In serial systems, a current source selects electrode pairs for current injection, and voltage measurement modules sequentially acquire boundary voltages channel-by-channel. In parallel systems, the current source similarly selects electrode pairs for current injection, and multiple voltage measurement modules simultaneously acquire boundary voltages from several channels, significantly enhancing measurement efficiency.
EIT image reconstruction is inherently an ill-posed inverse problem [2] due to having far fewer measurement points than the number of pixels in the reconstructed image. As a result, its spatial resolution remains significantly lower than that of traditional imaging techniques such as CT and MRI. However, EIT’s advantages include being non-invasive, radiation-free, and capable of real-time imaging, offering significant clinical potential.
EIT is widely utilized in clinical fields such as lung function monitoring [3, 4], brain function monitoring [5], cancer detection [6], and human-machine interfaces [7]. Lung function monitoring is particularly prevalent, as EIT provides real-time visualizations of pulmonary gas distribution, assisting clinicians in optimizing ventilator settings, thus reducing risks associated with ventilator-induced lung injury.
Commercial EIT systems include Sentec LuMon [8], Dräger Pulmovista 500 [9], Timpel Enlight 2100 [10], Gense [11], and Sciospec’s EIT16 and EIT32/64/128+ systems [12]. Notably, the Dräger Pulmovista 500 system offers real-time visualization of pulmonary gas distribution, helping clinicians assess lung collapse, overinflation, and ventilation heterogeneity. Sciospec’s systems typically employ 16 electrodes but also support configurations up to 256 electrodes, achieving real-time frame rates of 100 fps with frequencies ranging from 100 Hz to 1 MHz.
Recent advancements in EIT technology include systems like Sheffield Mk3.5 [13], ACT-4 [14], ACT-5 [15], OXBACT-3 [16], OXBACT-5 [17], UCLH Mk2 [18], UCLH Mk2.5 [19], KHU Mark2 [20], KHU Mark2.5 [21, 22], mfEIT [23], and Dartmouth College-developed systems [24, 25, 26]. Among these, the Sheffield Mk3.5 system can measure impedance at 30 frequency points ranging from 2 kHz to 1.6 MHz, with an average signal-to-noise ratio (SNR) of 40 dB across all frequencies. This system utilizes 8 electrodes and includes 8 identical data acquisition boards, each equipped with a digital signal processor (DSP) to generate driving frequencies and perform fast Fourier transform (FFT) demodulation. The ACT-5 system, employing 32 electrodes, can display conductivity distribution images at a rate of 27 frames per second. Its adaptive current source can apply fully programmable current patterns over a frequency range of 5 kHz to 500 kHz. Additionally, it can simultaneously display electrocardiogram (ECG) readings during the imaging process. The KHU Mark2.5 system features automatic calibration, continuous operation at 250 kHz frequency for 24 hours, and maintains an SNR above 80 dB. This FPGA-based system includes multiple impedance measurement modules, each containing an independent constant current source, a differential voltage measurement module, and a current-source calibrator.
Although the electrical impedance tomography (EIT) devices mentioned above offer excellent imaging performance, their high costs limit their widespread adoption. As a result, developing a low-cost, easy-to-implement EIT system is crucial. In clinical medicine, EIT is primarily used for real-time lung function monitoring. Therefore, the device must be capable of capturing respiratory changes in the lungs while also being compact, portable, and capable of fast data acquisition to meet the demands of dynamic imaging.
This paper aims to develop a portable and low-cost EIT system based on the AD5933 chip. Firstly, according to the impedance characteristics of the human body, current injection is realized by applying an excitation voltage to the object under test and measuring the corresponding excitation current. Compared to traditional methods using a current source, this approach does not require the complete elimination of the DC offset inherent in current sources. Meanwhile, to tackle issues such as oscillations and leakage currents caused by parasitic capacitances in multiplexers, the system implements various optimization measures, enhancing stability at higher frequencies. Secondly, to improve data acquisition speed, the system employs five AD5933s operating in parallel. The TCA9548A is used to multiplex the I2C bus, while the ICS553 provides a unified external clock to all AD5933s, ensuring high synchronization during parallel acquisition. Finally, the PCB layout was optimized for a more compact design. Additionally, Bluetooth wireless communication and lithium battery power supply were integrated to further enhance the device’s portability. The structure of this paper is organized as follows: Section II introduces the developed EIT system and its specifications. Section III presents experimental results obtained from measurements on a water tank and a male subject. Section IV provides a brief summary.
2 Methods
2.1 System Architecture
The impedance hardware acquisition system is designed around the AD5933, which integrates a high-precision direct digital synthesizer (DDS) capable of generating sine wave signals within a frequency range of 1 kHz to 100 kHz. Additionally, the AD5933 incorporates a 12-bit analog-to-digital converter (ADC) for high-resolution analog signal acquisition. Most importantly, it includes a digital signal processor (DSP) that allows for accurate measurement of both the amplitude and phase of the sine wave signal.
The overall architecture of the EIT system is shown in Fig.1(a), and the physical prototype in Fig.1(b), with all circuit modules labeled. The system consists of five AD5933 units, a control module, a power supply module, a Bluetooth transmission module, a voltage amplification module, an I-V conversion module, five voltage measurement modules, an I2C bus control module, an external clock module, and multiple multiplexer modules. During the impedance acquisition process, the STM32 first controls the first AD5933 via the I2C bus to generate a sine voltage signal, which is then amplified into a sine voltage excitation by the voltage amplification module. The output of the voltage amplification module and the current input of the I-V conversion circuit are connected to two electrodes for voltage excitation through two multiplexer modules. Meanwhile, the AD5933 selects either the output terminal connected to the boundary voltage measurement or the output terminal of the I-V conversion circuit to acquire either the boundary voltage data or the current data generated by the sine voltage excitation passing through the object under test. The STM32 controls the selection of different pins on the multiplexer chip, enabling a complete scan of the object under test. The AD5933 transmits the processed voltage and current data to the STM32 via the I2C bus. Finally, the STM32 sends the calculated impedance data to the computer software for image reconstruction via a serial bus or Bluetooth.
The system offers high flexibility, allowing configuration into various measurement modes based on actual requirements. By adjusting the jumper settings, the enable (EN) pins of the multiplexers connected to the voltage detection circuitry can be set to either a high or low level, enabling or disabling multiplexer channels. This setup helps avoid signal conflicts and interference during measurements, while also reducing parasitic capacitance effects from unused multiplexers. It can operate in a two-terminal serial measurement mode for basic impedance detection, or switch to a four-terminal serial measurement mode, where current is injected through two electrodes and voltage is measured across the other two, thereby minimizing the effect of contact resistance on the results and enhancing accuracy. Additionally, the system supports a four-terminal semi-parallel measurement mode, enabling simultaneous voltage measurements across multiple electrodes, further accelerating the data acquisition process.
In the two-terminal series mode of measurement, the impedance is calculated based on the excitation current generated by the fixed excitation voltage passing through the test object. At this moment, the system architecture is shown in Fig.2.
Figure 2: EIT System Architecture in the Two-Terminal Serial Mode.
In the four-terminal series measurement mode, a single voltage excitation can only obtain the voltage of one electrode or the current generated by the excitation voltage passes through the object. Therefore, channel switching via multiplexers is necessary to measure all boundary voltages and excitation currents in the adjacent excitation mode. At this point, the system architecture is shown in Fig.3.
Figure 3: EIT System Architecture in the Four-Terminal Serial Mode.
In the four-terminal semi-parallel measurement mode, by using five AD5933 chips in parallel, the system can simultaneously obtain the voltage of four electrodes and the current generated by the excitation voltage passing through the object during a single current injection. Through multiple channel switching, all boundary voltages and excitation currents under the adjacent excitation mode can be measured. At this point, the system architecture is shown in Fig.4.
Figure 4: EIT System Architecture in the Four-Terminal Semi-Parallel Mode.
2.2 Voltage Excitation And Current Measurement
2.2.1 Circuit structure analysis
An excitation voltage is applied to the test object, and the resulting excitation current is measured. In this manner, the measured current represents the actual excitation current injected into the test object. In contrast to the method that uses a fixed-amplitude current signal for excitation, this approach does not depend on the high output impedance of the current source to maintain current amplitude stability, nor does it require ensuring high linearity between the current amplitude and the control signal to guarantee measurement accuracy. Additionally, since a fixed-amplitude voltage excitation is used, the circuit does not need to dynamically adjust the output current based on the impedance characteristics of the test object. Another advantage of the voltage excitation method is that even if there is a small amount of DC bias in the excitation voltage, the resulting DC bias current will be minimal due to the human body’s extremely high impedance to DC. Therefore, there is no need to completely eliminate the DC bias voltage. Moreover, the circuit design is simple and reliable, requiring only a 3.3V power supply for stable operation.
Figure 5: Schematic of the Voltage Excitation Circuit.
Fig.5 shows the voltage excitation circuit, and the output voltage of the circuit is:
$V_{0}(s)=\frac{V_{cc}}{2}-\frac{R_{2}}{R_{1}}\cdot\frac{\frac{R}{2}}{\frac{1}{j\omega C_{1}}+\frac{R}{2}}V_{i}(s)$ (1)
When $R_{2}=R_{1}$, the output signal $V_{0}(s)$ has the same AC amplitude as the input signal $V_{i}(s)$, and the DC bias of the output signal $V_{0}(s)$ is $\frac{Vcc}{2}$.
Figure 6: Schematic of the I-V Conversion Circuit.
Fig.6 shows the I-V conversion circuit, and the output voltage of the circuit is:
$V_{o}(s)=-\frac{R_{f}}{1+sR_{f}C_{f}}\cdot I(s)$ (2)
The output terminal of the voltage excitation circuit and the current input terminal of the I-V conversion circuit are connected to both ends of the test object. The output signal of the voltage excitation circuit has an AC amplitude of 1.98 $V_{p-p}$ and a DC bias level of $\frac{Vcc}{2}$. Meanwhile, the voltage at the inverting input terminal of the operational amplifier in the I-V conversion circuit is also $\frac{Vcc}{2}$. Therefore, the DC voltage across the test object is 0V, while the AC voltage amplitude is 1.98 $V_{p-p}$.
2.2.2 Circuit stability analysis
In impedance acquisition devices, the use of multiplexing chips and the physical characteristics of the PCB inevitably introduce parasitic capacitance at both the inverting input of the operational amplifier in the I-V conversion circuit and the output of the operational amplifier in the voltage excitation circuit. Parasitic capacitance can degrade circuit stability and increase the risk of oscillation. Shen Qidong (1990) studied the impact of parasitic capacitance on operational amplifier circuits and proposed effective compensation methods [27], which serve as valuable references for the circuit analysis in this study.
Figure 7: I-V Conversion Circuit with Parasitic Capacitance Effects.
As shown in Fig.7, the voltage excitation signal $V_{1}$ generates an excitation current through resistor $R_{1}$, which is applied to the inverting input of the operational amplifier in the I-V conversion circuit. Meanwhile, the parasitic capacitance introduced by the multiplexer, which exists between the inverting input of the operational amplifier and ground, influences the circuit’s frequency characteristics. The loop gain of the circuit is:
$A(s)\cdot F(s)=\frac{A_{0}}{1+\frac{s}{\omega_{0}}}\cdot\frac{R_{1}}{R_{1}+R_{f}}\frac{1}{1+\frac{s}{\omega_{P}}}$ (3)
In the formula, $A_{0}$ and ${\omega_{0}}$ denote the open-loop DC gain and the 3 dB bandwidth of the operational amplifier, respectively. Moreover, $\omega_{P}=\frac{1}{R_{P}\cdot C_{1}}$, $R_{p}=R_{1}\parallel R_{f}$.
In the feedback circuit, the combination of capacitor $C_{1}$ and resistors $R_{1}$ and $R_{f}$ introduces a lagging component. When $C_{1}$ or the feedback resistor $R_{f}$ is relatively large, the second pole frequency associated with the lagging component decreases significantly, making the circuit more susceptible to oscillation. To resolve this, a capacitor $C_{f}$ is typically connected in parallel with the feedback resistor $R_{f}$ for lead compensation. The function of $C_{f}$ is to introduce a zero, and when $C_{f}=\frac{R_{1}\cdot C_{1}}{R_{f}}$, the zero and the pole coincide, thereby compensating the system’s frequency characteristics and effectively preventing circuit instability and oscillation.
Figure 8: Voltage Excitation Circuit with Parasitic Capacitance Effects.
As shown in Fig.8, this represents the second-stage section of the voltage excitation circuit. The parasitic capacitance of the multiplexer serves as the capacitive load $C_{L}$ for the operational amplifier, forming a lagging component together with the op-amp’s internal output resistance $R_{0}$. This combination influences the circuit’s frequency characteristics.
The closed-loop gain of the circuit is:
$A_{F}(s)=\frac{A(s)}{1+A(s)\cdot F(s)}=\frac{\frac{A_{0}}{1+\frac{s}{\omega_{0}}}\frac{1}{1+\frac{s}{\omega_{Q}}}}{1+\frac{A_{0}}{1+\frac{s}{\omega_{0}}}\frac{1}{1+\frac{s}{\omega_{Q}}}\cdot\frac{R_{1}}{R_{1}+R_{f}}}$ (4)
Here, ${\omega_{0}}$ represents the intrinsic pole frequency of the operational amplifier, and $\omega_{Q}=\frac{1}{C_{L}\cdot R_{0}}$ denotes the second pole frequency introduced by the output capacitive load and the op-amp’s output resistance. The expression $A_{F}(s)$ is rewritten as:
$A_{F}(s)$ $=\frac{\frac{R_{1}+R_{f}}{R_{1}}}{1+\frac{R_{1}+R_{f}}{R_{1}A_{0}}\frac{s(\omega_{0}+\omega_{Q})}{\omega_{0}\omega_{Q}}+\frac{R_{1}+R_{f}}{R_{1}A_{0}}\frac{s^{2}}{\omega_{0}\omega_{Q}}}$
$=\frac{\frac{R_{1}+R_{f}}{R_{1}}}{1-\left(\frac{\omega}{\omega_{h}}\right)^{2}+2j\xi\frac{\omega}{\omega_{h}}}$ (5)
The damping coefficient $\xi$ of the circuit is expressed as:
$\xi=\frac{\omega_{0}+\omega_{Q}}{2\omega_{h}}=\sqrt{\frac{(\omega_{0}+\omega_{Q})^{2}}{{4}\cdot\frac{R_{1}A_{0}}{R_{1}+R_{f}}{\omega_{0}\omega_{Q}}}}$ (6)
where $\omega_{h}$ is: $\omega_{h}=\sqrt{\frac{R_{1}A_{0}}{R_{1}+R_{f}}\omega_{0}\omega_{Q}}$, Since $\omega_{0}\ll\omega_{Q}$, we approximate:
$\xi\approx\sqrt{\frac{1}{4\cdot\frac{R_{1}A_{0}}{R_{1}+R_{f}}w_{0}C_{L}R_{0}}}$ (7)
From the expression for the damping coefficient, $\xi$ is inversely proportional to the output resistance $R_{0}$ of the operational amplifier. A lower output resistance $R_{0}$ enhances the damping effect, effectively suppressing system oscillations. The output impedance of the AD8606 increases with both frequency and gain. At 100 kHz with a gain of 100, its output impedance remains below 20 $\Omega$, ensuring compliance with system requirements.
2.3 Multiplexer and Voltage Measurement
Channel switching is implemented using the ADG706 multiplexer. The ADG706 is controlled via four binary address lines (A0, A1, A2, and A3) to select one of the 16 input terminals (S1–S16) and connect it to the common output terminal D. Compared with traditional relay-based switching methods, the use of the ADG706 may introduce parasitic capacitance effects. Nevertheless, it is still selected as the channel switching device due to its smaller size and faster switching speed.
In EIT, a constant-amplitude current signal is commonly used for excitation. In this context, the impedance of the region under test directly influences the boundary voltage. Specifically, a higher impedance results in a larger boundary voltage, while a lower impedance leads to a smaller one. To ensure that the boundary voltage remains within the ADC’s sampling range, a programmable gain amplifier (PGA) is typically employed to adjust the signal and prevent it from exceeding the input limits. However, when the boundary voltage is very low, the PGA must apply a high gain, which may also amplify noise, thereby degrading the system’s signal-to-noise ratio (SNR). In this study, the adjacent excitation-adjacent measurement mode is adopted, using a constant-amplitude AC voltage to excite two adjacent electrodes among the 16 electrodes. In this mode, the maximum boundary voltage difference between the two adjacent electrodes is equal to the amplitude of the excitation voltage. Since the boundary voltage difference has a relatively large amplitude and remains within the ADC’s maximum input range, it can be ensured that the boundary voltage stays within a suitable range for sampling. Therefore, with fixed-amplitude voltage excitation, a PGA is no longer required to adjust the boundary voltage. Instead, voltage acquisition can be accomplished using only a voltage follower circuit.
When using a four-terminal semi-parallel configuration for impedance acquisition and selecting two electrodes for voltage excitation, the four voltage detection circuits are connected to four electrodes via four 16-to-1 multiplexer chips. In this configuration, if the 16 input terminals (S1-S16) of each multiplexer chip are directly connected to the 16 electrodes, each voltage detection electrode introduces a parasitic capacitance of 200 pF. This capacitance results from the activation of the input channels (S1-S16) on the multiplexer chips. When the signal frequency exceeds 50 kHz, parasitic capacitance may induce leakage currents, which can distort the current field distribution within the object under test and compromise measurement accuracy. To mitigate this issue, the 16 input terminals of the four multiplexer chips should not be connected directly to the electrodes but instead routed through voltage followers (i.e. buffers), as shown in Fig.9. Owing to their high input impedance and low output impedance, voltage followers effectively suppress leakage currents, thereby ensuring accurate current field distribution and precise measurement results.
Figure 9: Electrode Signal Acquisition and Multiplexer Circuit.
2.4 I2C Multiplexer and External Clock
This paper implements communication with five AD5933 chips through the I2C multiplexer chip TCA9548A. Developed by Texas Instruments (TI), the TCA9548A is capable of splitting a single I2C bus signal into eight independent channels, each of which can be individually enabled or disabled. This approach allows multiple I2C devices with identical addresses to share the same I2C bus, effectively avoiding address conflicts and enabling simultaneous activation of several I2C channels. By appropriately configuring the registers of the TCA9548A, simultaneous transmission of measurement commands to five AD5933 chips was successfully realized, ensuring synchronized data acquisition.
This paper selects the ICS553 as the clock distribution chip. Developed by Renesas, the ICS553 is a clock buffer capable of generating up to four output clock signals from a single reference clock input. Employing an external clock primarily ensures synchronized data acquisition. The ADC sampling frequency of the AD5933 is determined by dividing the external clock frequency by 16. Thus, using a unified external clock allows the ADC sampling frequencies of the five parallel AD5933 chips to remain consistent. Fig.10(a) illustrates the connection between the clock signal outputs of the ICS553 and the clock signal inputs of the AD5933. The purpose of this circuit is to limit the signal amplitude. The equivalent circuit is presented in Fig.10(b).
Figure 10: (a)Schematic of Clock Signal Connection from ICS553 to AD5933. (b)Equivalent Circuit Model.
Let $V_{1}$ be the clock voltage signal output from pins $Q_{0}$ to $Q_{3}$ of the ICS553, and $V_{2}$ be the clock voltage signal received at the MCLK pin of the AD5933. The calculation formula for $V_{2}$ is as follows:
$V_{2}=\frac{\text{VCC}}{2}\cdot\frac{R_{0}}{R_{0}+R}+V_{1}\cdot\frac{R}{R_{0}+R}$ (8)
2.5 Microcontroller Program Flow
2.5.1 Serial Measurement
After the system is powered on, the STM32 writes specific values to the control register of the TCA9548A to connect to the first AD5933 and initialize it.
In the two-terminal serial impedance measurement process, the multiplexer selects two electrodes from the 16 available electrodes to apply a voltage excitation, while the first AD5933 measures the current. In the four-terminal serial impedance measurement process, the multiplexer selects two electrodes from the 16 available electrodes to perform a total of 17 voltage excitations. During the first excitation, the first AD5933 measures the current; from the second excitation onward, the first AD5933 sequentially measures the voltage of each electrode until all 16 electrodes have been measured.
During current or voltage measurements, STM32 sends commands to the control register of the AD5933 to put it into standby mode, initialize the start frequency, and begin frequency scanning. STM32 then continuously queries the status register of the AD5933 to confirm whether the Discrete Fourier Transform (DFT) has been completed. If completed, STM32 reads the real and imaginary part data from the AD5933 registers.
The voltage excitation electrodes are then changed via the multiplexer. After each change in the two electrodes used for voltage excitation, STM32 performs an impedance calculation. Once all impedance data collection is completed, STM32 transmits a complete frame of impedance data to the PC for further processing. Fig.11(a) shows the electrode excitation and measurement sequence in the two-terminal serial measurement mode. Fig.11(b) shows the excitation and measurement sequence in the four-terminal serial measurement mode. Fig.13(a) illustrates the program flowchart for both the two-terminal and four-terminal serial measurement modes.
Figure 11: (a)Two-Terminal Serial Scanning Sequence. (b)Four-Terminal Serial Scanning Sequence. (c)Four-Terminal Parallel Scanning Sequence.
2.5.2 Parallel Measurement
After the system is powered on, STM32 first writes specific values to the control register of the TCA9548A to connect the first, second, third, fourth, and fifth AD5933 devices and complete their initialization.
In the four-terminal half-parallel impedance measurement process, the multiplexer selects two electrodes from the 16 available electrodes to apply a total of 5 voltage excitations. During each excitation, the first AD5933 measures the current, while the second, third, fourth, and fifth AD5933 devices measure the voltage of one electrode from the 16 electrodes. During the first voltage excitation, the second, third, fourth, and fifth AD5933 devices measure the voltage at electrodes 1, 5, 9, and 13, respectively; during the second excitation, they measure the voltage at electrodes 2, 6, 10, and 14, respectively; during the third excitation, they measure the voltage at electrodes 3, 7, 11, and 15, respectively; during the fourth excitation, they measure the voltage at electrodes 4, 8, 12, and 16, respectively; and during the fifth excitation, they measure the voltage at electrodes 5, 9, 13, and 1, respectively. This ensures that the voltage differences between adjacent electrodes (e.g., 1-2, 2-3, 3-4, 4-5) are measured by the second AD5933, the voltage differences between other adjacent electrodes (e.g., 5-6, 6-7, 7-8, 8-9) are measured by the third AD5933, and so on. This design ensures that the voltage difference between each pair of adjacent electrodes is measured by the same AD5933, which improves the common-mode rejection ratio of the system’s voltage measurements.
During the current or voltage measurement process, STM32 first writes specific values to the TCA9548A register to connect it with the first AD5933, which is responsible for generating the voltage excitation signal. Then, STM32 sends a command to the control register of the first AD5933 to put it into standby mode, initialize the start frequency, and begin frequency scanning. This allows the AD5933 to generate the excitation voltage and measure the excitation current after the voltage passes through the test specimen. Next, STM32 writes new values to the TCA9548A register to connect it with the second, third, fourth, and fifth AD5933 devices, and follows the same steps by sending standby commands, initializing start frequency commands, and starting frequency scanning commands to these AD5933 for voltage measurement. During the Discrete Fourier Transform (DFT), the excitation signal is applied first, followed by the measurement of boundary voltages. At the same time, the frequency of the excitation signal is set to an integer multiple of the DFT data acquisition frequency for 1024 points. This ensures that each 1024-point acquisition window contains a complete number of excitation signal cycles, avoiding measurement errors caused by incomplete waveforms. A specific diagram is shown in Fig.12. The formula for DFT data sampling frequency is:
$f_{DFT}=\frac{f_{ADC}}{n}$ (9)
Where $f_{ADC}$ is the ADC sampling frequency and n is the number of collected samples (such as 1024 samples). It can be calculated that when the clock frequency is 16 MHz, the ADC sampling frequency is 1 MHz, and the DFT data sampling frequency is $\frac{1\,\text{MHz}}{1024}$.
Figure 12: (a)Sample Set Spanning Complete Integer Excitation Periods. (b)Sample Set Not Spanning Complete Integer Excitation Periods.
Once the frequency scan begins, STM32 writes specific values to the TCA9548A register to sequentially connect it with the first, second, third, fourth, and fifth AD5933 devices. With each connection, STM32 continuously queries the status register of the AD5933 to confirm whether the Discrete Fourier Transform (DFT) has been completed. If completed, STM32 reads the real and imaginary part data from the AD5933 registers.
The voltage excitation electrodes are then switched by multiplexing. After each switch of the two excitation electrodes, the system calculates the impedance based on one current measurement and 16 voltage measurements from the electrodes, resulting in 16 impedance values. After collecting 256 impedance data points from 16 electrode pairs in the adjacent excitation-adjacent measurement mode, STM32 transmits a complete frame of impedance data to the PC for further processing. Fig.11(c) shows the excitation and measurement sequence in the four-terminal semi-parallel measurement mode. Fig.13(b) illustrates the program flowchart of the four-terminal semi-parallel impedance measurement system.
Figure 13: (a)Serial Measurement Method Program Flowchart. (b)Parallel Measurement Method Program Flowchart.
2.6 Impedance Calculation
When using the two-terminal measurement method, it is important to understand that the measured data does not directly represent the real and imaginary components of the impedance. Instead, these data reflect the frequency-domain representation of the current signal. Since the excitation voltage is known, the real and imaginary components of the impedance can be calculated using the measured current data in conjunction with a predetermined gain factor.
The phase angle is given by the following formula:
$\phi=\tan^{-1}\left(\frac{Re}{Im}\right)$ (10)
Here, $Re$ represents the real part, and $Im$ represents the imaginary part. The phase angle of the current signal measured by the AD5933 results from various phase shifts introduced by different elements in the measurement system. These include the phase of the excitation voltage generated by the AD5933, the phase shift of the current signal relative to the voltage signal within the sample under test, the phase shift introduced by the I-V conversion circuitry, and the phase shift from the AD5933’s internal low-pass filter. These combined phase shifts contribute to the final measured phase angle.
To measure the phase shift of the current signal relative to the voltage signal in the sample under test, a two-step approach is required. Step 1: Connect a pure resistor to the circuit and measure the phase angle of the resulting signal with the AD5933. The measured phase at this point represents the system phase, which does not include the phase shift of the current signal relative to the voltage signal within the sample. Instead, it includes the excitation voltage phase generated by the AD5933, the phase shift introduced by the I-V conversion circuitry, and the phase shift from the AD5933’s internal low-pass filter. Step 2: Replace the pure resistor with the actual sample under test and measure again. Now, the phase shift of the current signal relative to the voltage signal in the sample can be determined using the following formula: $\phi_{cal}=\phi_{unk}-\phi_{res}$. Here, $\phi_{res}$ is the system phase measured in Step 1 with the known pure resistor, and $\phi_{unk}$ is the total phase measured in Step 2 with the actual sample. By calculating the difference between these two phase measurements, the phase shifts introduced by the system itself can be eliminated, allowing for the calculation of the phase shift of the current signal relative to the voltage signal within the sample.
If: $|I|\cdot e^{j\phi_{unk}}=Re_{unk}+Im_{unk}j$, $|I|\cdot e^{j\phi_{res}}=Re_{res}+Im_{res}j$, where $Re_{unk}$ and $Im_{unk}$ represent the real and imaginary parts of the current signal measured by the AD5933 when the test sample is connected, and $Re_{res}$ and $Im_{res}$ represent the real and imaginary parts of the current signal measured when a pure resistor is connected. According to the phase correction formula $\phi_{cal}=\phi_{unk}-\phi_{res}$, we obtain:
$|I|\cdot e^{j\phi_{cal}}=\left(Re_{unk}+Im_{unk}j\right)\frac{\left(Re_{res}-Im_{res}j\right)}{\sqrt{Re_{res}^{2}+Im_{res}^{2}}}$ (11)
Next, define the complex expression for the calibrated current signal as: $|I|\cdot e^{j\phi_{cal}}=Re_{cal}+Im_{cal}j$, we then have:
$Re_{cal}+Im_{cal}j=\left(Re_{unk}+Im_{unk}j\right)\frac{\left(Re_{res}-Im_{res}j\right)}{\sqrt{Re_{res}^{2}+Im_{res}^{2}}}$ (12)
By solving Eq.12, the following results:
$\begin{cases}Re_{cal}=\frac{Re_{unk}Re_{res}+Im_{unk}Im_{res}}{\sqrt{Re_{res}^{2}+Im_{res}^{2}}}\\ Im_{cal}=\frac{Re_{res}Im_{unk}-Re_{unk}Im_{res}}{\sqrt{Re_{res}^{2}+Im_{res}^{2}}}\end{cases}$ (13)
For a two-terminal system, the complex impedance Z can be calculated as:
$Z=\frac{1}{{Gain}\cdot|I|\cdot e^{j\phi_{cal}}}=\frac{1}{{Gain}\cdot\left(Re_{cal}+Im_{cal}j\right)}$ (14)
Here, $Gain$ is the gain factor, a constant determined during the calibration process. The magnitude of the impedance $|Z|$ is calculated by:
$|Z|=\frac{1}{Gain\cdot\sqrt{Re_{cal}^{2}+Im_{cal}^{2}}}$ (15)
The imaginary part of the impedance, $Z_{Im}$, is calculated as:
$Z_{Im}=\frac{-Im_{cal}}{\text{Gain}\cdot\left(Re_{cal}^{2}+Im_{cal}^{2}\right)}$ (16)
When using the four-terminal measurement method, the voltage signal from one of the 16 electrodes is processed using the Discrete Fourier Transform (DFT), resulting in the following expression: $X_{1}(f)=Re_{elec1}+Im_{elec1}j$, This result is calculated by:
$X_{1}(f)=\textstyle\sum_{n=0}^{N-1}x_{1}(n)\cdot e^{-j\frac{2\pi}{N}kn},\quad k=0,1,2,\dots,N-1$ (17)
where $X_{1}(f)$ is the frequency-domain component at frequency k, $x_{1}(n)$ is the discrete-time signal acquired by the ADC, and $e^{-j\frac{2\pi}{N}kn}$ is the rotation factor. $N$ is the total number of points used in the DFT calculation. DFT satisfies linearity, which means for any constants $a$ and $b$, and any signals $x(n)$ and $y(n)$, the following holds:
${DFT}\{ax(n)+by(n)\}=aX(f)+bY(f)$ (18)
Thus, for the voltage difference signal between two electrodes, $x_{1}(n)-x_{2}(n)$, its DFT can be expressed as:
${DFT}\{x_{1}(n)-x_{2}(n)\}=X_{1}(f)-X_{2}(f)$ (19)
By substituting the DFT results of the voltage signals, $X_{1}(f)=Re_{elec1}+Im_{elec1}j$ and $X_{2}(f)=Re_{elec2}+Im_{elec2}j$, into the above equation, we obtain:
${DFT}\{x_{1}(n)-x_{2}(n)\}$ $=\left(Re_{\text{elec1}}-Re_{\text{elec2}}\right)$
$+\left(Im_{\text{elec1}}-Im_{\text{elec2}}\right)j$ (20)
Therefore, the DFT of the voltage difference between two electrodes can be obtained by individually performing DFT on each electrode’s voltage signal and then calculating the differences of their real and imaginary parts.
If the DFT result of the voltage difference signal between the two electrodes is: $X_{diff}(f)=\left(Re_{elec1}-Re_{elec2}\right)+\left(Im_{elec1}-Im_{elec2}\right)j$, and the DFT result of the current signal is: $X_{I}(f)=Re_{I}+Im_{I}j$, the complex impedance $Z(f)$ can be calculated as:
$Z(f)={Gain}\cdot\frac{X_{\text{diff}}(f)}{X_{I}(f)}$ (21)
Similarly, $Gain$ is the gain factor determined during calibration. The magnitude of the complex impedance $|Z(f)|$ is calculated as:
$|Z(f)|=\sqrt{Z_{Re}(f)^{2}+Z_{Im}(f)^{2}}$ (22)
2.7 Graphical User Interface
PyQt5 allows developers to build graphical user interfaces (GUIs) using Python. In this article, we will use PyQt5 to design and implement the GUI. The system software is composed of six key components: data reading thread, data saving thread, calibration data thread, imaging thread, interface mapping thread, and the main program. With multi-threading support, tasks like real-time data processing and image updates are handled by separate threads, ensuring a smooth and responsive main interface without lag or delays.
Fig.14 illustrates the GUI. This interface not only supports real-time imaging and data visualization but also allows for data saving and calibration. While users analyze the EIT images, they can conveniently save the data for later offline analysis and experiment reproduction. On the left side, a dynamic line chart displays the impedance data collected by the acquisition device. The horizontal axis represents 256 impedance points, while the vertical axis shows their corresponding values. The real-time updating chart provides a clear visualization of the impedance data distribution, helping users monitor changes and detect potential anomalies. On the right side, the EIT image is displayed. The colors in the image correspond to different impedance values, with red indicating low impedance areas and blue indicating high impedance areas. This color differentiation helps users more clearly observe the impedance distribution within the scanned area. At the bottom of the interface, various functional buttons are provided, including ”Start,” ”Pause,” and ”Calibration.” Users can effortlessly start or stop data acquisition, refresh the interface, save data, and control imaging with these buttons. The layout of the buttons is clean and intuitive, greatly enhancing user convenience.
Figure 14: Screenshot of the User Interface During Data Collection.
3 Results
3.1 System Performance Evaluation
To verify the performance of the impedance acquisition device, the collected data was initially reviewed. In a water tank with a bottom diameter of 20 cm, 3 cm of water with a conductivity of 150 $\mu\text{S}/\text{cm}$ was added, and no objects were placed inside the tank. The impedance acquisition device was used to collect impedance data under two measurement modes: four-terminal serial and four-terminal semi-parallel, with a frequency range set from 8 kHz to 100 kHz. To avoid frequency leakage, when using a 16 MHz external crystal oscillator, the frequency was set to integer multiples of 1 MHz/1024. Fig.15 shows the data collected at 9766 Hz and 97656 Hz under the two measurement modes.
From the data acquisition results, it can be observed that the data quality was high across both measurement modes, which preliminarily verifies that the device supports a working frequency up to 97656 Hz.
Figure 15: Comparison of Impedance Data across Measurement Modes at 9766 Hz and 97656 Hz.
Subsequently, under the same experimental conditions, the system’s signal-to-noise ratio (SNR) and relative standard deviation (RSD) were measured in the water tank. For each frequency, 100 measurements were taken to calculate the SNR and RSD. The SNR was calculated using Eq.23 [25].
$SNR=10\log_{10}\left(\frac{\sum_{i=1}^{N}Z_{i}^{2}}{\sum_{i=1}^{N}(Z_{i}-\bar{Z})^{2}}\right)$ (23)
Where $Z_{i}$ represents the value of each measurement, $\bar{Z}$ is their average, and N is the total number of measurements (in this case, N = 100). The $\sum_{i=1}^{N}Z_{i}^{2}$ in the formula represents the total signal strength, and $\sum_{i=1}^{N}(Z_{i}-\bar{Z})^{2}$ represents the noise component. The RSD was calculated using Eq.24 [25].
$RSD=\frac{\sqrt{\frac{1}{N}\sum_{i=1}^{N}(Z_{i}-\bar{Z})^{2}}}{\bar{Z}}\times 100\%$ (24)
Where $Z_{i}$ represents the value of each measurement, $\bar{Z}$ is their average, and N is the total number of measurements (again, N = 100). The standard deviation $\sqrt{\frac{1}{N}\sum_{i=1}^{N}(Z_{i}-\bar{Z})^{2}}$ of the measurement data, divided by the mean $\bar{Z}$, indicates the degree of dispersion or variability of the data.
Fig.16(a) and Fig.16(b) show the SNR and RSD at various frequencies, respectively. In both the four-terminal serial and four-terminal semi-parallel measurement modes, the SNR values remain between 50 dB and 70 dB, while the RSD values stay below 0.4% across all frequencies. This indicates that the measurement data exhibit small fluctuations and high stability.
Figure 16: (a)SNR vs Frequency in Parallel and Series Modes. (b)RSD vs Frequency in Parallel and Series Modes.
Finally, under the same experimental conditions, the system’s reciprocity error (RE) was measured in the water tank using adjacent excitation and adjacent measurement modes. The RE was calculated using Eq.25 [28].
$RE=\frac{v(c,d)\mid I(a,b)-v(a,b)\mid I(c,d)}{v(c,d)\mid I(a,b)}$ (25)
Where $v(c,d)\mid I(a,b)$ represents the voltage difference measured between electrodes c and d when current is applied between electrodes a and b, and $v(a,b)\mid I(c,d)$ represents the voltage difference measured between electrodes a and b when current is applied between electrodes c and d.
Fig.17 shows the RE at various frequencies. At all frequencies, the RE values are below 2%, indicating that the system exhibits good reciprocity.
Figure 17: RE vs Frequency in Parallel and Series Modes.
3.2 Water Tank Phantom Imaging Results
This study employs the difference imaging method, which requires two frames of data: one as reference data and the other as target data. Difference imaging consists of time-difference imaging and frequency-difference imaging. Time-difference imaging is based on impedance data collected at different time points, while frequency-difference imaging is based on impedance data collected at different frequency points. In subsequent experiments, sweet potatoes and plastic bottles of various shapes and sizes were placed in water, and impedance magnitude data collected by the impedance acquisition device were used for time-difference imaging experiments. Image reconstruction was carried out using the one-step Gauss-Newton method and implemented in the EIDORS toolbox of MATLAB R2020b (MathWorks, USA). The experimental results demonstrate the imaging performance of the system in time-difference amplitude imaging mode under the two-terminal serial, four-terminal serial, and four-terminal semi-parallel measurement modes, and compare the performance across different frequencies. The specific imaging results are shown in Fig.18.
Figure 18: Comparison of Time-Difference Amplitude Images across Modes and Frequencies.
In the imaging map, the blue areas represent high impedance, while the red areas represent low impedance. According to the experimental results, the imaging color of the sweet potato changes with the test frequency: at lower frequencies, the sweet potato appears blue, while at higher frequencies, it appears red. This is because the impedance of biological tissues decreases as the frequency increases. At low frequencies, the impedance of the sweet potato is higher than that of water, while at higher frequencies, it is lower than that of water. Furthermore, the impedance difference between the sweet potato and water is more pronounced at both low and high frequencies, resulting in clearer imaging. In contrast, the plastic bottle in the tank always appears blue and remains clear across all frequencies, as its impedance does not change with frequency and is consistently higher than that of water.
In subsequent experiments, frequency-difference imaging was performed using impedance magnitude data collected by the impedance acquisition device, with the sweet potato submerged in water. The same device was then used to collect data for frequency-difference imaging tests with submerged potatoes and plastic bottles. The experimental results demonstrate the system’s imaging performance in frequency-difference amplitude imaging mode under the two-terminal serial, four-terminal serial, and four-terminal semi-parallel measurement modes, and compare the performance of reference data and target data at different frequencies. The specific imaging results are shown in Fig.19.
Figure 19: Comparison of Frequency-Difference Amplitude Images across Modes and Frequencies.
The imaging results show that when a sweet potato is placed in water for frequency-difference imaging, the region corresponding to it appears red, indicating a lower impedance value. This is because the impedance of the sweet potato decreases with increasing frequency. Similarly, when a potato and plastic bottle are placed in water for frequency-difference imaging, the region corresponding to the potato also appears red, while the region corresponding to the plastic bottle shows no color. This is because the impedance of the potato decreases with increasing frequency, while the impedance of the plastic bottle remains unchanged and unaffected by frequency. Furthermore, when the reference data frequency is lower and the target data frequency is higher (i.e., when the frequency difference is larger), the imaging of the sweet potato and potato is clearer. This is because a larger frequency difference results in a more significant impedance variation in the sweet potato and potato. Moreover, compared to the two-terminal measurement method, the four-terminal method more accurately reflects the size and shape of the sweet potato. This is because the four-terminal method includes more impedance information in each data frame and effectively suppresses contact impedance interference.
Next, time-difference imaging experiments using the imaginary component of impedance data were conducted with the submerged sweet potato, followed by similar experiments with the submerged plastic bottle. The experimental results demonstrate the imaging performance of the device in time-difference imaginary part imaging mode under the two-terminal serial, four-terminal serial, and four-terminal semi-parallel measurement modes, and compare the device’s performance at different frequencies. The specific imaging results are shown in Fig.20.
Figure 20: Comparison of Time-Difference Imaginary-Part Images across Modes and Frequencies.
The imaging results show that the region corresponding to the carrot appears red, while the region corresponding to the plastic bottle appears blue. This is because the carrot, as a biological entity made up of cells, exhibits capacitive characteristics, while the plastic bottle, as a non-biological material, shows different impedance characteristics. The significance of imaginary part imaging lies in its ability to reveal the phase information of impedance data, in contrast to amplitude imaging. In the biomedical field, due to differences in the conductivity and capacitive properties of biological tissues and other materials, imaginary part imaging provides additional spatial information, which helps distinguish between different types of tissues or materials.
3.3 Calf Cross-Section Imaging Results
In the experiment to reconstruct the cross-sectional impedance image of human calf tissue, data collection was conducted on a 28-year-old male subject with a Body Mass Index (BMI) of 20.5. Sixteen copper electrodes were evenly arranged around the calf using a PCB electrode band and connected to the impedance acquisition device, as shown in the Fig.21(a). The impedance acquisition device was then used to collect the impedance magnitude data of the calf tissue at two frequencies, 8789 Hz and 97656 Hz, under the four-terminal serial measurement mode. Finally, frequency-difference imaging was applied to reconstruct the cross-sectional image of the calf. Image reconstruction was performed using the one-step Gauss-Newton method, implemented in the EIDORS toolbox of MATLAB R2020b (MathWorks, USA). Fig.22 shows the comparison of the impedance frequency-difference imaging results of the calf cross-section with the anatomical structures.
The experimental results show that frequency-difference imaging can effectively reflect the tissue structure of the calf cross-section. For example, the tibia and fibula regions appear blue, while the blood vessels and nerve regions appear red. This is because different tissues (such as muscle, fat, blood, and bone) exhibit varying conductivity changes at different frequencies. Blood vessels and nerves, which contain more water and electrolytes, show significant impedance variation with frequency, while the impedance of the fibula and tibia changes less with frequency. These results demonstrate the broad potential of EIT for non-invasive detection and tissue identification.
Figure 21: (a)Arrangement of 16 Electrodes Around the Calf Using an Electrode Belt. (b)Arrangement of 16 Electrodes Around the Thorax Using an Electrode Belt.
Figure 22: Comparison of Calf Frequency-Difference Imaging with Its Anatomical Structure.
3.4 Pulmonary Respiration Imaging Results
In the experiment of reconstructing dynamic lung breath images, data collection was similarly conducted on a 28-year-old male subject, with a measurement frequency of 97656 Hz. Sixteen copper electrodes were evenly arranged around the thorax using a PCB electrode band and connected to the impedance acquisition device, as shown in the Fig.21(b). Next, the subject was instructed to hold their breath in a resting state, and the device collected impedance data as reference data. The subject then resumed normal breathing, and the device continued to collect impedance data as target data. Finally, time-difference imaging was used to reconstruct the dynamic changes in lung breathing. The reconstruction process used one-step Gauss-Newton method. Fig.23 and Fig.24 show the imaging results of lung breathing changes under four-terminal serial and four-terminal semi-parallel measurement modes.
The imaging maps include five complete breathing cycles, with each row representing one full breathing cycle and each column displaying images from different stages of breathing. In the imaging maps, the blue areas represent high-impedance regions, while the red areas represent low-impedance regions. From these images, it is clear that as breathing progresses, the impedance values in different lung regions change. During the inhalation phase, gas enters the chest, and the impedance of the corresponding lung region increases, appearing as blue. During the exhalation phase, gas exits the chest, and the impedance of the corresponding lung region decreases, appearing as red. The results indicate that the impedance acquisition device has excellent accuracy and stability, and is capable of effectively capturing the dynamic changes in lung breathing.
Figure 23: Pulmonary Respiration Imaging with Four-Terminal Serial Mode.
Figure 24: Pulmonary Respiration Imaging with Four-Terminal Semi-Parallel Mode.
4 Conclusion
To address the issues of high cost and bulkiness in existing electrical impedance tomography (EIT) devices, this study presents a high-precision, high-speed, low-cost, and portable impedance acquisition device. The main research contents include:
  1. 1.
    Development of an impedance acquisition device capable of flexibly switching among three measurement modes: two-terminal serial, four-terminal serial, and four-terminal parallel, depending on application requirements. Instead of employing a current source for excitation, the device utilizes voltage excitation and measures the resulting excitation current passing through the body. This approach eliminates the need for dynamic adjustment of excitation current based on impedance variations and avoids the necessity of fully eliminating DC offset, resulting in enhanced stability. To enhance portability and acquisition speed, the system employs multiplexers instead of relays for channel switching. However, the introduction of multiplexers brings parasitic capacitance issues. Therefore, several measures were implemented to suppress associated signal oscillations during voltage excitation and current measurement and to minimize leakage currents during voltage measurement. To improve acquisition speed, the device utilizes five AD5933 chips operating in parallel, with I2C bus control managed via the TCA9548 and a unified external clock provided by the ICS553, ensuring synchronous multi-channel data acquisition. Additionally, portability is enhanced by using lithium battery power and Bluetooth-based wireless data transmission.
  2. 2.
    Development of a graphical user interface (GUI) using PyQt5, streamlining user operations and facilitating real-time imaging and data acquisition.
  3. 3.
    Performance evaluation of the impedance acquisition device. Initial experiments using a water tank phantom demonstrated effective imaging capabilities across various measurement modes (two-terminal and four-terminal, serial and parallel) and imaging modes (time-difference and frequency-difference) up to 100 kHz. Subsequent tests involved successful dynamic lung respiration imaging using the time-difference method, effectively capturing respiratory movements. Finally, frequency-difference imaging was utilized for cross-sectional imaging of calf tissue, successfully distinguishing anatomical features such as bones, blood vessels, and nerves.