### v2

# Design of a Wearable Parallel Electrical Impedance Imaging System for Healthcare

Abstract
A wireless wearable Electrical Impedance Tomography (EIT) system has been developed utilizing the AD5933 chip to achieve real-time imaging of lung respiration. The system employs a voltage excitation method tailored to human impedance characteristics, injecting current by applying a known voltage and measuring the resulting current through the body. Additionally, specific measures have been implemented to effectively suppress signal oscillations and leakage currents caused by parasitic capacitances. To enhance data acquisition speed, the system employs five parallel AD5933 units, with multiple techniques implemented to ensure high synchronization during simultaneous measurements. Performance testing shows that the system achieves a signal-to-noise ratio greater than 50 dB, a relative standard deviation below 0.3%, and a reciprocity error under 0.8%. Imaging experiments using a water tank phantom, human lungs during breathing, and a resting human calf further demonstrate that this portable EIT system can accurately measure biological tissues with high precision and low cost.
{IEEEkeywords} Electrical Impedance Tomography, Parallel Architecture, AD5933, Portable Measurement System
1 Introduction
[…]
Electrical impedance tomography (EIT) [1] is a non-invasive imaging technique in which a low-amplitude, safe alternating current is injected into the target region through electrode pairs, while the resulting boundary voltages induced by the current are measured at the remaining electrodes. The internal impedance distribution of the target region is then reconstructed using inverse-problem algorithms.
Figure 1: (a)Overview of the EIT System Architecture. (b)Photograph of the EIT System.
There are primarily two methods for measuring boundary voltages. In serial systems, a current source selects electrode pairs for current injection, and voltage measurement modules sequentially acquire boundary voltages channel-by-channel. In parallel systems, the current source similarly selects electrode pairs for current injection, and multiple voltage measurement modules simultaneously acquire boundary voltages from several channels, significantly enhancing measurement efficiency.
EIT image reconstruction is inherently an ill-posed inverse problem [2] due to having far fewer measurement points than the number of pixels in the reconstructed image. As a result, its spatial resolution remains significantly lower than that of traditional imaging techniques such as CT and MRI. However, EIT’s advantages include being non-invasive, radiation-free, and capable of real-time imaging, offering significant clinical potential.
EIT is widely used in clinical fields such as lung monitoring [3, 4], brain assessment [5], cancer detection [6], and human-machine interfaces [7]. Among these, lung monitoring is the most common, as EIT provides real-time visualization of pulmonary gas distribution, helping clinicians optimize ventilator settings and reduce the risk of lung injury.
Commercial EIT systems include Sentec LuMon [8], Dräger Pulmovista 500 [9], Timpel Enlight 2100 [10], Gense [11], and Sciospec’s EIT16 and EIT32/64/128+ systems [12]. Notably, the Dräger Pulmovista 500 system provides real-time visualization of pulmonary gas distribution, aiding clinicians in assessing lung collapse, overinflation, and ventilation heterogeneity. Sciospec’s EIT systems typically use 16 electrodes but support configurations up to 256, offering real-time frame rates up to 100 fps and operating across a frequency range of 100 Hz to 1 MHz.
Recent advancements in EIT technology include systems like Sheffield Mk3.5 [13], ACT-4 [14], ACT-5 [15], OXBACT-3 [16], OXBACT-5 [17], UCLH Mk2 [18], UCLH Mk2.5 [19], KHU Mark2 [20], KHU Mark2.5 [21, 22], mfEIT [23], and Dartmouth College-developed systems [24, 25, 26]. Among these, the Sheffield Mk3.5 system can measure impedance at 30 frequency points ranging from 2 kHz to 1.6 MHz, with an average signal-to-noise ratio (SNR) of 40 dB across all frequencies. This system utilizes 8 electrodes and includes 8 identical data acquisition boards, each equipped with a digital signal processor (DSP) to generate driving frequencies and perform fast Fourier transform (FFT) demodulation. The ACT-5 system, equipped with 32 electrodes, can generate conductivity distribution images at a rate of 27 frames per second. Its adaptive current source supports fully programmable current patterns across a frequency range of 5 kHz to 500 kHz, and it can simultaneously display ECG readings during imaging. The KHU Mark2.5 system features automatic calibration and supports continuous operation at a frequency of 250 kHz for up to 24 hours, while maintaining a SNR above 80 dB. Based on FPGA architecture, the system incorporates multiple impedance measurement modules, each equipped with an independent constant current source, a differential voltage measurement unit, and a current source calibrator.
Although the EIT devices mentioned above deliver excellent imaging performance, their high cost limits widespread adoption. Therefore, developing a low-cost, easy-to-implement EIT system is essential. In clinical medicine, EIT is primarily used for real-time monitoring of lung function. Therefore, the device must be compact and portable, with rapid data acquisition capabilities to effectively capture dynamic respiratory changes in the lungs.
This paper aims to develop a portable and low-cost EIT system based on the AD5933 chip. Firstly, according to the impedance characteristics of the human body, current injection is realized by applying an excitation voltage to the object under test and measuring the corresponding excitation current. Compared to traditional methods using a current source, this approach does not require the complete elimination of the DC offset inherent in current sources. Meanwhile, to tackle issues such as oscillations and leakage currents caused by parasitic capacitances in multiplexers, the system implements various optimization measures, enhancing stability at higher frequencies. Secondly, to improve data acquisition speed, the system employs five AD5933 chips operating in parallel, with I2C bus control managed via the TCA9548A and a unified external clock provided by the ICS553, ensuring synchronized data acquisition across all AD5933 channels. Finally, the PCB layout was optimized for a more compact structure. Additionally, Bluetooth wireless communication and lithium battery power were adopted to further improve the device’s portability.
The structure of this paper is organized as follows: Section II introduces the developed EIT system and its specifications. Section III presents experimental results obtained from measurements on a water tank and a male subject. Section IV provides a brief summary.
2 Methods
2.1 System Architecture
The impedance acquisition system is built around the AD5933, which features an integrated direct digital synthesizer (DDS) capable of generating sine wave signals over a frequency range of 1 kHz to 100 kHz. It also includes a 12-bit analog-to-digital converter (ADC) and a digital signal processor (DSP), allowing for precise measurement of both the amplitude and phase of the sine wave signals.
Fig.1(a) illustrates the overall architecture of the EIT system, while Fig.1(b) shows the physical prototype with all circuit modules clearly labeled. The system consists of an STM32 development board, a power supply module, a Bluetooth transmission module, a voltage excitation module, a buffer module, a data acquisition module, an I2C multiplexer module, an external clock module, and a multiplexer module.
During the impedance acquisition process, the STM32 first uses the I2C bus to control the first AD5933 to generate a sine wave voltage signal. This signal is then converted into an excitation voltage by the voltage excitation module. The output of the voltage excitation circuit and the input of the current measurement circuit are connected to a pair of electrodes via multiplexers to apply voltage excitation. Meanwhile, the AD5933 selects either the output of the voltage measurement circuit or that of the current measurement circuit to acquire the boundary voltage or the current generated by the voltage excitation passing through the object under test. The STM32 controls the channel selection of the multiplexer to perform a full scan of the object under test. The AD5933 transmits the processed voltage and current data to the STM32 via the I2C bus. Finally, the STM32 sends the computed impedance data to a computer via Bluetooth for image reconstruction.
The system is highly flexible and can be configured into different measurement modes based on actual requirements. By adjusting the jumper settings, the enable (EN) pins of the multiplexers connected to the voltage detection circuitry can be set to high or low levels to enable or disable the corresponding multiplexers, thereby reducing the parasitic capacitance caused by unused multiplexers. It can be configured in a two-terminal serial mode for basic impedance analysis, or in a four-terminal serial mode, which separates the current injection and voltage sensing paths to minimize the impact of contact resistance and improve measurement accuracy. Additionally, the system supports a four-terminal parallel mode, enabling simultaneous voltage measurements on multiple electrodes to greatly accelerate data acquisition.
In the two-terminal series measurement mode, impedance is calculated based on the current generated by applying a fixed voltage excitation across the test object. The system architecture for this mode is illustrated in Fig.2.
Figure 2: EIT System Architecture in the Two-Terminal Serial Mode.
In the four-terminal series measurement mode, a single voltage excitation allows measurement of either the voltage at one electrode or the current resulting from the voltage excitation passing through the object. Therefore, channel switching via multiplexers is required to measure all boundary voltages and currents under the adjacent excitation configuration. The system architecture for this mode is shown in Fig.3.
Figure 3: EIT System Architecture in the Four-Terminal Serial Mode.
In the four-terminal parallel measurement mode, five AD5933 chips are used in parallel, enabling the system to simultaneously acquire the voltages of four electrodes and the current resulting from the voltage excitation passing through the object during a single voltage excitation. By performing multiple channel switches, the system can measure all boundary voltages and currents under the adjacent excitation configuration. The system architecture is shown in Fig.4.
Figure 4: EIT System Architecture in the Four-Terminal Parallel Mode.
2.2 Voltage Excitation And Current Measurement
2.2.1 Circuit structure analysis
An excitation voltage is applied to the test object, and the resulting excitation current is measured. In this way, the measured current represents the actual excitation current injected into the test object. Compared to the approach that uses a fixed-amplitude current signal for excitation, this method does not rely on a high output impedance of the current source to stabilize the current amplitude, nor does it require a highly linear relationship between the current amplitude and the control signal to ensure measurement accuracy. Moreover, by employing a fixed-amplitude voltage excitation, the circuit eliminates the need to dynamically adjust the output current in response to the impedance characteristics of the test object. Another benefit of employing voltage excitation is that a small DC bias in the excitation signal results in only a negligible DC current, owing to the human body’s very high impedance to direct current. Consequently, complete elimination of the DC bias voltage is not required. Furthermore, the circuit is both simple and robust, and can operate reliably with a 3.3V power supply.
Figure 5: Schematic of the Voltage Excitation Circuit.
Fig.5 illustrates the voltage excitation circuit. The output voltage is expressed as follows:
$V_{0}(s)=\frac{V_{cc}}{2}-\frac{R_{2}}{R_{1}}\cdot\frac{\frac{R}{2}}{\frac{1}{j\omega C_{1}}+\frac{R}{2}}V_{i}(s)$ (1)
Figure 6: Schematic of the I-V Conversion Circuit.
Fig.6 illustrates the I-V conversion circuit. The output voltage is expressed as follows:
$V_{o}(s)=-\frac{R_{f}}{1+sR_{f}C_{f}}\cdot I(s)$ (2)
The output of the voltage excitation circuit and the current input of the I-V conversion circuit are connected across the test object. The voltage excitation circuit outputs a 1.98 $V_{p-p}$ AC signal superimposed on a DC bias of $\frac{Vcc}{2}$. Since the input of the I-V conversion circuit is also held at $\frac{Vcc}{2}$, the DC voltage across the test object is effectively 0 V, with an AC amplitude of 1.98 $V_{p-p}$.
2.2.2 Circuit stability analysis
In impedance acquisition devices, the use of multiplexers and the physical characteristics of the PCB inevitably lead to parasitic capacitance at both the input of the I-V conversion circuit and the output of the voltage excitation circuit. Parasitic capacitance can compromise circuit stability and increase the risk of oscillation. Shen Qidong (1990) investigated the effects of parasitic capacitance on operational amplifier circuits and proposed effective compensation methods [27], which provide important references for the circuit analysis in this study.
Figure 7: I-V Conversion Circuit with Parasitic Capacitance Effects.
As shown in Fig.7, the voltage excitation signal $V_{1}$ generates a current through resistor $R_{1}$, which is applied to the input of the I-V conversion circuit. At the same time, the parasitic capacitance from the multiplexer, located between the inverting input of the operational amplifier and ground, affects the circuit’s frequency response. The loop gain of the circuit is:
$A(s)\cdot F(s)=\frac{A_{0}}{1+\frac{s}{\omega_{0}}}\cdot\frac{R_{1}}{R_{1}+R_{f}}\frac{1}{1+\frac{s}{\omega_{P}}}$ (3)
In the formula, $A_{0}$ and ${\omega_{0}}$ denote the low-frequency open-loop gain and the dominant pole frequency of the operational amplifier, respectively. Moreover, $\omega_{P}=\frac{1}{R_{P}\cdot C_{1}}$, $R_{p}=R_{1}\parallel R_{f}$.
In a feedback circuit, the combination of capacitor $C_{1}$ and resistors $R_{1}$ and $R_{f}$ forms a lagging network. When either $C_{1}$ or the feedback resistor $R_{f}$ is large, the second pole frequency introduced by this network drops significantly, making the circuit more prone to oscillation. To address this issue, a capacitor $C_{f}$ is commonly placed in parallel with the feedback resistor $R_{f}$ to provide lead compensation. $C_{f}$ introduces a zero, and when $C_{f}=\frac{R_{1}\cdot C_{1}}{R_{f}}$, this zero coincides with the pole, effectively compensating the system’s frequency response and helping to prevent instability and oscillation.
Figure 8: Voltage Excitation Circuit with Parasitic Capacitance Effects.
As shown in Fig.8, this is the second-stage section of the voltage excitation circuit. The parasitic capacitance from the multiplexer acts as a capacitive load $C_{L}$ on the operational amplifier. Together with the op-amp’s internal output resistance $R_{0}$, it forms a lagging network that affects the circuit’s frequency response.
The closed-loop gain of the circuit is:
$A_{F}(s)=\frac{A(s)}{1+A(s)\cdot F(s)}=\frac{\frac{A_{0}}{1+\frac{s}{\omega_{0}}}\frac{1}{1+\frac{s}{\omega_{Q}}}}{1+\frac{A_{0}}{1+\frac{s}{\omega_{0}}}\frac{1}{1+\frac{s}{\omega_{Q}}}\cdot\frac{R_{1}}{R_{1}+R_{f}}}$ (4)
Here, ${\omega_{0}}$ represents the dominant pole frequency of the operational amplifier, $\omega_{Q}=\frac{1}{C_{L}\cdot R_{0}}$, ${\omega_{Q}}$ denotes the second pole frequency introduced by the capacitive load and the op-amp’s output resistance. The expression $A_{F}(s)$ can be rewritten as:
$A_{F}(s)$ $=\frac{\frac{R_{1}+R_{f}}{R_{1}}}{1+\frac{R_{1}+R_{f}}{R_{1}A_{0}}\frac{s(\omega_{0}+\omega_{Q})}{\omega_{0}\omega_{Q}}+\frac{R_{1}+R_{f}}{R_{1}A_{0}}\frac{s^{2}}{\omega_{0}\omega_{Q}}}$
$=\frac{\frac{R_{1}+R_{f}}{R_{1}}}{1-\left(\frac{\omega}{\omega_{h}}\right)^{2}+2j\xi\frac{\omega}{\omega_{h}}}$ (5)
The damping factor $\xi$ of the circuit is:
$\xi=\frac{\omega_{0}+\omega_{Q}}{2\omega_{h}}=\sqrt{\frac{(\omega_{0}+\omega_{Q})^{2}}{{4}\cdot\frac{R_{1}A_{0}}{R_{1}+R_{f}}{\omega_{0}\omega_{Q}}}}$ (6)
where $\omega_{h}=\sqrt{\frac{R_{1}A_{0}}{R_{1}+R_{f}}\omega_{0}\omega_{Q}}$. Since $\omega_{0}\ll\omega_{Q}$, we have:
$\xi\approx\sqrt{\frac{1}{4\cdot\frac{R_{1}A_{0}}{R_{1}+R_{f}}w_{0}C_{L}R_{0}}}$ (7)
From the expression of the damping coefficient, it is evident that $\xi$ is inversely proportional to the output resistance $R_{0}$ of the operational amplifier. A lower $R_{0}$ increases the damping effect, thereby helping to suppress system oscillations. The output impedance of the AD8606 rises with both frequency and gain. Even at a frequency of 100 kHz and a gain of 100, the AD8606 maintains an output impedance below 20 $\Omega$, which meets the system’s requirements.
2.3 Multiplexer and Voltage Measurement
Channel switching utilizes the ADG706 multiplexer. Although the ADG706 can introduce parasitic capacitance compared to traditional relay-based methods, it remains the preferred option due to its smaller size and faster switching speed.
In EIT, excitation is commonly performed using a constant-amplitude current signal. In this setup, the impedance of the region under test directly influences the resulting boundary voltage. To ensure that this boundary voltage remains within the sampling range of the ADC, a Programmable Gain Amplifier (PGA) is typically employed to dynamically adjust the voltage level. However, in this study, excitation is performed using a constant-amplitude voltage signal. As a result, a PGA is no longer necessary for adjusting the boundary voltage, and voltage acquisition can be achieved using only a voltage follower circuit.
In the four-terminal parallel mode, if the four voltage detection circuits are directly connected to four electrodes via four multiplexer chips, each of these electrodes introduces a parasitic capacitance of 200 pF. This parasitic capacitance results from the activation of the multiplexer channels. At frequencies above 50 kHz, leakage currents induced by parasitic capacitance may alter the current field distribution within the object under test, leading to reduced measurement accuracy. To prevent this, the 16 input terminals of the four multiplexer chips should be isolated from the electrodes using voltage followers (buffers), as depicted in Fig.9.
Figure 9: Electrode Signal Acquisition and Multiplexer Circuit.
2.4 I2C Multiplexer and External Clock
Communication with five AD5933 chips is implemented through the I2C multiplexer chip TCA9548A. Developed by Texas Instruments (TI), the TCA9548A is capable of splitting a single I2C bus signal into eight independent channels, each of which can be individually enabled or disabled. This approach allows multiple I2C devices with identical addresses to share the same I2C bus, effectively avoiding address conflicts and enabling simultaneous activation of several I2C channels. By appropriately configuring the registers of the TCA9548A, simultaneous transmission of measurement commands to five AD5933 chips was successfully realized, ensuring synchronized data acquisition.
The ICS553 is selected as the clock distribution chip. Developed by Renesas, the ICS553 is a clock buffer capable of generating up to four output clock signals from a single reference clock input. Employing an external clock primarily ensures synchronized data acquisition. The ADC sampling frequency of the AD5933 is determined by dividing the external clock frequency by 16. Thus, using a unified external clock allows the ADC sampling frequencies of the five parallel AD5933 chips to remain consistent. Fig.10(a) illustrates the connection between the clock signal outputs of the ICS553 and the clock signal inputs of the AD5933. The purpose of this circuit is to limit the signal amplitude. The equivalent circuit is presented in Fig.10(b).
Figure 10: (a)Schematic of Clock Signal Connection from ICS553 to AD5933. (b)Equivalent Circuit Model.
Let $V_{1}$ be the clock voltage signal output from pins $Q_{0}$ to $Q_{3}$ of the ICS553, and $V_{2}$ be the clock voltage signal received at the MCLK pin of the AD5933. The calculation formula for $V_{2}$ is as follows:
$V_{2}=\frac{\text{VCC}}{2}\cdot\frac{R_{0}}{R_{0}+R}+V_{1}\cdot\frac{R}{R_{0}+R}$ (8)
2.5 Microcontroller Program Flow
2.5.1 Serial Measurement
After the system is powered on, the first AD5933 is connected and initialized. In the two-terminal serial impedance measurement process, the multiplexer selects two electrodes from the 16 available electrodes to apply a voltage excitation, while the first AD5933 measures the current. In the four-terminal serial impedance measurement process, the multiplexer selects two electrodes from the 16 available electrodes to perform a total of 17 voltage excitations. During the first excitation, the first AD5933 measures the current; from the second excitation onward, the first AD5933 sequentially measures the voltage of each electrode until all 16 electrodes have been measured.
During current or voltage measurements, first, a command to start frequency scanning is sent to the control register of the AD5933. Then, the status register of the AD5933 is continuously queried to check whether the Discrete Fourier Transform (DFT) has been completed. Once completed, the STM32 reads the real and imaginary part data from the AD5933 registers.
The voltage excitation electrodes are then changed via the multiplexer. After each change in the two electrodes used for voltage excitation, STM32 performs an impedance calculation. Once all impedance data collection is completed, STM32 transmits a complete frame of impedance data to the PC for further processing. Fig.12(a) shows the excitation and measurement sequence in the two-terminal serial measurement mode. Fig.12(b) shows the excitation and measurement sequence in the four-terminal serial measurement mode. Fig.13(a) illustrates the program flowchart for both the two-terminal and four-terminal serial measurement modes.
2.5.2 Parallel Measurement
After the system is powered on, each of the five AD5933 chips is connected and initialized. In the four-terminal parallel impedance measurement process, the multiplexer selects two electrodes from the 16 available electrodes to perform a total of five voltage excitations. During each excitation, the first AD5933 consistently measures the current, while the second to fifth AD5933 chips sequentially measure the voltages at specific electrodes as follows: in the first excitation, electrodes 1, 5, 9, and 13 are measured; in the second excitation, electrodes 2, 6, 10, and 14; in the third excitation, electrodes 3, 7, 11, and 15; in the fourth excitation, electrodes 4, 8, 12, and 16; and in the fifth excitation, electrodes 5, 9, 13, and 1. This design ensures that the voltage difference between each pair of adjacent electrodes is measured by the same AD5933, thereby improving the common-mode rejection ratio of the system’s voltage measurements.
During the current or voltage measurement process, the system first connects to the first AD5933 and sends a frequency scanning start command, causing it to generate an excitation voltage and measure the current. Subsequently, the system simultaneously connects to the second, third, fourth, and fifth AD5933 chips and sends a frequency scanning start command, enabling these AD5933 chips to perform voltage measurements. The method involves first applying the voltage excitation signal, followed by measuring the boundary voltages. Meanwhile, the frequency of the excitation signal is set to be an integer multiple of the data acquisition frequency for the 1024-point DFT. This ensures that each 1024-point acquisition window captures an integer number of excitation cycles, thereby avoiding measurement errors caused by incomplete waveforms. A detailed illustration is provided in in Fig.11. The formula for the DFT data sampling frequency is:
$f_{DFT}=\frac{f_{ADC}}{n}$ (9)
Where $f_{ADC}$ is the ADC sampling frequency and n is the number of collected samples (such as 1024 samples).
Figure 11: (a)Sample Set Spanning Complete Integer Excitation Periods. (b)Sample Set Not Spanning Complete Integer Excitation Periods.
Once the frequency scan begins, STM32 sequentially connects to the first, second, third, fourth, and fifth AD5933 chips. During each connection, it continuously queries the status register of the AD5933 to check whether the DFT has been completed. If completed, STM32 reads the real and imaginary part data from the AD5933 registers.
The voltage excitation electrodes are then changed via the multiplexer. After each change in the two electrodes used for voltage excitation, STM32 performs an impedance calculation. Once all impedance data collection is completed, STM32 transmits a complete frame of impedance data to the PC for further processing. Fig.12(c) shows the excitation and measurement sequence in the four-terminal parallel measurement mode. Fig.13(b) illustrates the program flowchart of the four-terminal parallel impedance measurement mode.
Figure 12: (a)Two-Terminal Serial Scanning Sequence. (b)Four-Terminal Serial Scanning Sequence. (c)Four-Terminal Parallel Scanning Sequence.
Figure 13: (a)Serial Measurement Method Program Flowchart. (b)Parallel Measurement Method Program Flowchart.
2.6 Impedance Calculation
In the two-terminal measurement mode, the measured quantity is the current. Given that the excitation voltage is fixed, the real and imaginary parts of the impedance can be derived from the current data together with a known gain factor.
To measure the phase shift of the current signal relative to the voltage signal in the sample under test, a two-step process is required. In the first step, a pure resistor is connected to the circuit for measurement. The phase angle measured by the AD5933 at this point represents the system phase. In the second step, the pure resistor is replaced with the sample under test, and another measurement is taken. The phase angle measured by the AD5933 now represents the total phase. The phase shift of the current signal relative to the voltage signal in the sample can be calculated using the formula: $\phi_{cal}=\phi_{unk}-\phi_{res}$, where $\phi_{res}$ is the system phase and $\phi_{unk}$ is the total phase. By subtracting the two values, the phase shift introduced by the system itself is eliminated, allowing the phase shift of the current signal relative to the voltage signal in the sample to be determined.
If: $|I|\cdot e^{j\phi_{unk}}=Re_{unk}+Im_{unk}j$ and $|I|\cdot e^{j\phi_{res}}=Re_{res}+Im_{res}j$, where $Re_{unk}$ and $Im_{unk}$ represent the real and imaginary parts of the current signal measured by the AD5933 when the test sample is connected, and $Re_{res}$ and $Im_{res}$ represent the real and imaginary parts of the current signal measured when a pure resistor is connected. According to the phase correction formula $\phi_{cal}=\phi_{unk}-\phi_{res}$, the complex representation of the calibrated current signal is:
$|I|\cdot e^{j\phi_{cal}}=\left(Re_{unk}+Im_{unk}j\right)\frac{\left(Re_{res}-Im_{res}j\right)}{\sqrt{Re_{res}^{2}+Im_{res}^{2}}}$ (10)
the complex impedance Z can be calculated as:
$Z=\frac{1}{{Gain}\cdot|I|\cdot e^{j\phi_{cal}}}=\frac{1}{{Gain}\cdot\left(Re_{cal}+Im_{cal}j\right)}$ (11)
$Gain$ is the gain factor, a constant determined during the calibration process.
In the four-terminal measurement mode, the voltage signals $x_{1}(n)$ and $x_{2}(n)$ from two adjacent electrodes are transformed using the DFT. The results are: $X_{1}(f)=Re_{elec1}+Im_{elec1}j$, $X_{2}(f)=Re_{elec2}+Im_{elec2}j$. The DFT satisfies the linearity property, so the DFT result of the voltage difference signal between the two electrodes is:
${DFT}\{x_{1}(n)-x_{2}(n)\}$ $=X_{1}(f)-X_{2}(f)$
$=\left(Re_{\text{elec1}}-Re_{\text{elec2}}\right)$
$+\left(Im_{\text{elec1}}-Im_{\text{elec2}}\right)j$ (12)
If the DFT result of the current signal is: $X_{I}(f)=Re_{I}+Im_{I}j$, the complex impedance $Z(f)$ can be calculated as:
$Z(f)={Gain}\cdot\frac{X_{1}(f)-X_{2}(f)}{X_{I}(f)}$ (13)
2.7 Graphical User Interface
PyQt5 enables developers to build graphical user interfaces (GUIs) using Python. In this work, we use PyQt5 to design and implement the GUI. The system software consists of six main components: the data reading thread, data saving thread, calibration thread, imaging thread, interface mapping thread, and the main program. By introducing multi-threading, operations such as real-time data processing and image updates can be performed in background threads, ensuring the main interface remains smooth and responsive without stuttering or lag.
Fig.14 shows the GUI, which integrates real-time imaging and data visualization, as well as data saving and calibration functions. While intuitively analyzing the EIT images, users can also easily save the data for subsequent offline analysis and experimental reproduction.
On the left side, a dynamic line chart presents the impedance data collected by the acquisition device. The horizontal axis represents 256 impedance points, while the vertical axis shows their corresponding values. This real-time updating chart offers an intuitive view of the impedance distribution, enabling users to track changes and identify potential anomalies. On the right side, the EIT image is displayed. The image uses color to represent different impedance values—red indicates regions of low impedance, while blue indicates high impedance. This visual differentiation helps users clearly observe the impedance distribution within the scanned area. At the bottom of the interface, multiple functional buttons are available, such as ”Start,” ”Pause,” and ”Calibration.” These buttons allow users to easily start or stop data acquisition, refresh the interface, save data, and manage imaging settings. The layout is clean and user-friendly, significantly improving operational efficiency and ease of use.
Figure 14: Screenshot of the User Interface During Data Collection.
3 Results
3.1 System Performance Evaluation
To evaluate the performance of the impedance acquisition device, a preliminary inspection of the collected data was conducted. A water tank with a bottom diameter of 20 cm was filled with water to a depth of approximately 3 cm. The water had a conductivity of 150 $\mu\text{S}/\text{cm}$, and no objects were placed in the tank during the measurement. Impedance data were collected using the impedance acquisition device under two measurement modes: four-terminal serial and four-terminal parallel. The frequency range was set from 8 kHz to 100 kHz, and to minimize frequency leakage, the excitation frequencies within this range were configured as integer multiples of 1 MHz/1024 when using a 16 MHz crystal oscillator. Fig.15 presents the data collected at frequencies of 9766 Hz and 97656 Hz under both measurement modes.
From the data acquisition results, both measurement modes demonstrated high data standardization, providing preliminary validation that the device can operate reliably at frequencies up to 97656 Hz.
Figure 15: Comparison of Impedance Data across Measurement Modes at 9766 Hz and 97656 Hz.
Subsequently, under the same experimental conditions, the system’s signal-to-noise ratio (SNR) and relative standard deviation (RSD) were measured in the water tank. At each frequency, 100 measurements were conducted to calculate the SNR and RSD. The SNR was calculated using Eq.14, while the RSD was determined according to Eq.15 [25].
$SNR=10\log_{10}\left(\frac{\sum_{i=1}^{N}Z_{i}^{2}}{\sum_{i=1}^{N}(Z_{i}-\bar{Z})^{2}}\right)$ (14)
$RSD=\frac{\sqrt{\frac{1}{N}\sum_{i=1}^{N}(Z_{i}-\bar{Z})^{2}}}{\bar{Z}}\times 100\%$ (15)
Here, $Z_{i}$ denotes the value of each individual measurement, $\bar{Z}$ represents the average of these values, and N is the number of measurements (in this case, N = 100).
Fig.16(a) and Fig.16(b) show the SNR and RSD measured at various frequencies, respectively. In both the four-terminal serial and four-terminal parallel measurement modes, the SNR remains within the range of 50 to 70 dB, while the RSD stays below 0.3% across all frequencies. These results indicate minimal fluctuation and high stability in the measurement data.
Figure 16: (a)SNR vs Frequency in Parallel and Series Modes. (b)RSD vs Frequency in Parallel and Series Modes.
Finally, under the same experimental conditions, the system’s reciprocity error (RE) was measured in the water tank using the adjacent excitation and adjacent measurement modes. The RE was calculated according to Eq.16 [28].
$RE=\frac{v(c,d)\mid I(a,b)-v(a,b)\mid I(c,d)}{v(c,d)\mid I(a,b)}$ (16)
Here, $v(c,d)\mid I(a,b)$ denotes the voltage difference measured between electrodes c and d when current is applied between electrodes a and b, while $v(a,b)\mid I(c,d)$ denotes the voltage difference measured between electrodes a and b when current is applied between electrodes c and d.
Fig.17 shows the RE measured at various frequencies. At all frequencies, the RE remains below 0.8%, demonstrating that the system possesses good reciprocity.
Figure 17: RE vs Frequency in Parallel and Series Modes.
3.2 Water Tank Phantom Imaging Results
The difference imaging method requires two frames of data: one serving as the reference and the other as the target. There are two types of difference imaging: time-difference imaging and frequency-difference imaging. Time-difference imaging utilizes impedance data acquired at different time points, whereas frequency-difference imaging uses data collected at different frequencies.
In subsequent experiments, impedance magnitude data collected by the impedance acquisition device were used to perform time-difference imaging on sweet potatoes and plastic bottles of various shapes and sizes submerged in water. Fig.18 presents the imaging results. The imaging reconstruction was performed using the one-step Gauss-Newton method, implemented with the EIDORS toolbox in MATLAB R2020b (MathWorks, USA).
According to the experimental results, the imaging color of the sweet potato changes with the test frequency: it appears blue at lower frequencies and turns red at higher frequencies. This is due to the frequency-dependent impedance of biological tissues, which decreases as the frequency increases. At low frequencies, the impedance of the sweet potato is higher than that of water, whereas at high frequencies, it becomes lower. Moreover, the impedance contrast between the sweet potato and water is more pronounced at both low and high frequencies, leading to clearer imaging results. In contrast, the plastic bottle in the water tank consistently appears blue across all frequencies, as its impedance remains constant and is significantly higher than that of water.
Additionally, impedance magnitude data from the same acquisition device were employed for frequency-difference imaging, initially on sweet potatoes submerged in water and subsequently on potatoes and plastic bottles. Fig.19 illustrates the outcome of the imaging.
The imaging results demonstrate that when a sweet potato is placed in water for frequency-difference imaging, its corresponding region appears red, indicating a lower impedance value. This occurs because the impedance of the sweet potato decreases as the frequency increases. Similarly, when a potato and a plastic bottle are placed in the water, the region corresponding to the potato also appears red, whereas the plastic bottle’s region remains uncolored. This is due to the fact that the potato’s impedance likewise decreases with increasing frequency, while the impedance of the plastic bottle remains constant and is unaffected by frequency changes.
Finally, using the imaginary part of impedance data gathered by the same device, time-difference imaging was conducted on submerged carrots, followed by similar imaging on plastic bottles. Fig.20 demonstrates the imaging results.
The imaging results indicate that the region corresponding to the carrot appears red, while the area corresponding to the plastic bottle appears blue. This distinction arises because the carrot, as a biological tissue composed of cells, exhibits capacitive properties, whereas the plastic bottle, being a non-biological material, displays distinct electrical impedance characteristics.
Figure 18: Time-Difference Imaging Result Based on Impedance Amplitude Data.
Figure 19: Frequency-Difference Imaging Result Based on Impedance Amplitude Data.
Figure 20: Time-Difference Imaging Result Based on Imaginary Part of Impedance Data.
3.3 Calf Cross-Section Imaging Results
In the experiment to reconstruct the cross-sectional impedance image of human calf tissue, data collection was conducted on a 28-year-old male subject with a Body Mass Index (BMI) of 20.5. Firstly, Sixteen copper electrodes were evenly arranged around the calf using a PCB-based electrode band and connected to the impedance acquisition device, as shown in Fig.21(a). The device then measured the impedance magnitude of the calf tissue at two frequencies—8,789 Hz and 97,656 Hz—using the four-terminal serial measurement mode. Finally, frequency-difference imaging was employed to reconstruct the cross-sectional image of the calf. The reconstruction was carried out using the one-step Gauss-Newton method, implemented in the EIDORS toolbox within MATLAB R2020b (MathWorks, USA). Fig.22 presents a comparison between the impedance frequency-difference imaging results and the anatomical structures of the calf cross-section.
The experimental results demonstrate that frequency-difference imaging can effectively reveal the tissue structure of the calf cross-section. For example, the tibia and fibula regions appear blue, while blood vessels and nerve regions appear red. This contrast arises because blood vessels and nerves contain higher levels of water and electrolytes, resulting in more pronounced impedance variations with frequency. In contrast, the impedance of the tibia and fibula remains relatively stable across different frequencies.
Figure 21: (a)Arrangement of 16 Electrodes Around the Calf Using an Electrode Belt. (b)Arrangement of 16 Electrodes Around the Thorax Using an Electrode Belt.
Figure 22: Comparison of Calf Frequency-Difference Imaging with Its Anatomical Structure.
3.4 Pulmonary Respiration Imaging Results
In the experiment of reconstructing dynamic lung breath images, data collection was similarly conducted on a 28-year-old male subject, with a measurement frequency of 97656 Hz. Firstly, Sixteen copper electrodes were evenly arranged around the thorax using a PCB-based electrode band and connected to the impedance acquisition device, as shown in Fig.21(b). The subject was first instructed to hold their breath in a resting state, during which impedance data were collected as reference data. The subject then resumed normal breathing, and the device continued recording impedance data as target data. Finally, time-difference imaging was used to reconstruct the dynamic changes in lung respiration. The reconstruction was performed using the one-step Gauss-Newton method. Fig.23 and Fig.24 show the imaging results of pulmonary ventilation changes under the four-terminal serial and four-terminal parallel measurement modes.
The imaging maps display five complete breathing cycles, with each row corresponding to a full cycle and each column representing different phases of respiration. These images clearly illustrate the dynamic changes in impedance across various lung regions throughout the respiratory process. During inhalation, as air enters the thoracic cavity, the impedance in the relevant lung areas increases, appearing in blue. Conversely, during exhalation, as air leaves the thoracic cavity, the impedance decreases, which is visualized in red.
Figure 23: Pulmonary Respiration Imaging with Four-Terminal Serial Mode.
Figure 24: Pulmonary Respiration Imaging with Four-Terminal Parallel Mode.
4 Conclusion
To address the limitations of existing EIT devices, such as high cost and bulky design, this study proposes a portable impedance acquisition device characterized by high precision, fast measurement speed, and affordability. The primary research areas include:
  1. 1.
    Development of an impedance acquisition device capable of flexibly switching among three measurement modes: two-terminal serial, four-terminal serial, and four-terminal parallel, depending on application requirements. Instead of employing a current source for excitation, the device utilizes voltage excitation and measures the resulting excitation current passing through the body. This approach eliminates the need for dynamic adjustment of excitation current based on impedance variations and avoids the necessity of fully eliminating DC offset, resulting in enhanced stability. To enhance portability and acquisition speed, the system employs multiplexers instead of relays for channel switching. However, the introduction of multiplexers brings parasitic capacitance issues. Therefore, several measures were implemented to suppress associated signal oscillations during voltage excitation and current measurement and to minimize leakage currents during voltage measurement. To improve acquisition speed, the device utilizes five AD5933 chips operating in parallel, with I2C bus control managed via the TCA9548 and a unified external clock provided by the ICS553, ensuring synchronous multi-channel data acquisition. Additionally, portability is enhanced by using lithium battery power and Bluetooth-based wireless data transmission.
  2. 2.
    Developed a user-friendly GUI with PyQt5 to simplify operation procedures and support efficient real-time imaging and data collection.
  3. 3.
    Systematic performance evaluations were conducted on the EIT device, demonstrating excellent results: the SNR exceeded 50 dB, the RSD was below 0.3%, and the RE was less than 0.8%. To further assess its imaging capabilities, three experiments were carried out. First, a water tank phantom experiment showed that the device provided stable and accurate imaging at frequencies below 100 kHz. Second, time-difference imaging was used to visualize lung ventilation, successfully capturing the dynamic changes associated with the respiratory cycle. Third, frequency-difference imaging was applied to cross-sectional imaging of the calf, with the results indicating the preliminary differentiation of anatomical structures such as bones, blood vessels, and nerves.